A biochemistry laboratory student used the Bradford protein assay to measure the lysozyme (protein) content in egg whites. The Bradford protein assay is a spectrophotometric technique that uses a dye, Coomassie Brilliant Blue, whose absorption undergoes a spectroscopic shift when bound to a protein. In the unbound state, Coomassie Brilliant Blue displays a red color. As protein binds to the dye, its absorbance at 595 nm increases and the dye changes to a blue color. The student prepared a 3.900 mL sample containing 2 mg/mL Coomassie Brilliant Blue. To this sample, 500.0 μL of 75.0 μg/mL lysozyme was added and an absorbance of 0.245 was measured at 595 nm. Calculate the corrected absorbance for this sample.
Answer #1

Solution;

The computation are expressed below,

c=\frac{\frac{75}{500}}{3.90}*10^{-3}=0.03858*10^{-3} M

The molar extinct co-efficient is,

\varepsilon =(nW*5500)+(nY*1490)+(nC*125)

for bradford protein by plugging nW,nYand nC values we get,

\varepsilon =(0.5*5500)+(1*1490)+(1*125)=4365M^{-1}cm^{-1}

A=\varepsilon lc=4365*1*0.0385*10^{-3}=0.168

corrected absorbance = 0.245-0.168 =0.077